30 likes | 149 Vues
Following the recent earthquake, a thorough assessment of the MUMON equipment was conducted. Key findings include no visible damage to the walls or sensors in the MUMON room, and the moving stage operates normally with no damage to electric power cables. Electric power cables connecting NU3 and Muhut were damaged but have been repaired. Gas lines were checked for leaks and confirmed safe. While most electronics show no damage, additional checks on certain sensors and alignment are pending until power is fully restored in the Muhut facility.
E N D
MUMON (Suzuki) Moving stage MUMON room gas line 16 Si MUMON room and MUMON • Checked the outside and the inside by viewing • There seems to be no damage to the wall of the room and sensors • It was confirmed that the moving stage can be controlled from Muhut, and was not damaged. (Electric power cable of moving stand was not damaged) Signal and HV cables btw MUMON and Muhut • Checked by using test pulse and confirmed the reflection pulse by the oscilloscope
MUMON Electric power cable Gas-line 17 • Electric power cable btw NU3 and Muhut • Cut by the gap of the ground due to the earthquake • Repaired by facility group on Mar.12. • Gas-line • Confirmed: No leakage in whole gas line including MUMON chamber was found after initial leak check. (Further leak check will be done when all equipments are ready) • Some parts are distorted by the earthquake • The distorted part outside NU3 is tentatively removed for the repair work around NU3.
MUMON: Detail items 18 • PCs in Muhut: No damage • Electronics in Muhut: No damage -AVR is not turned on yet (Need Sakashita san’s check) • Equipments (incl. its cables) to measure the gas pressure, the temperature, the gas flow and the O2 which are located in NU3 - Pressure sensors were checked by tester, and not damaged. - O2 monitor was turned on, and worked OK. - Temperature and flow sensors will be check within few weeks. • Sensors (Si and IC) -Will be checked after the full recovery of the power in the Muhut. -The schedule is not decided yet • Alignment of MUMON - Need to be done. The schedule is not decided yet.